The damage threshold for the retina is 100 W/m2. This means that a damage to the retina can occur at an average intensity greater than this. The light coming from a 5.0 mW green laser has a diameter of 1.2 mm. (a) Will this laser exceed the damage threshold for the retina? (b) What are the maximum values of the electric and magnetic fields of the beam coming from this laser?


1
Expert's answer
2022-04-28T13:01:41-0400

"\dot W=\frac WS=4.4~\frac{kW}{m^2}>100~\frac W{m^2}=\dot W_0."